# Attractor Memory for Long-Term Time Series Forecasting: A Chaos Perspective

### v4

@@ Abstract
In long-term time series forecasting (LTSF) tasks, an increasing number of models have acknowledged that discrete time series originate from continuous dynamic systems and have attempted to model their dynamical structures. Recognizing the chaotic nature of real-world data, our model, Attraos, incorporates chaos theory into LTSF, perceiving real-world time series as observations from unknown high-dimensional chaotic dynamic systems. Under the concept of attractor invariance, Attraos utilizes non-parametric Phase Space Reconstruction embedding and the proposed multi-scale dynamic memory unit to memorize historical dynamics structure and predicts by a frequency-enhanced local evolution strategy. Detailed theoretical analysis and abundant empirical evidence consistently show that Attraos outperforms various LTSF methods on mainstream LTSF datasets and chaotic datasets with only one-twelfth of the parameters compared to PatchTST.
@@ 1 Introduction
In the intricate dance of time, time series unfold. Emerged from continuous dynamic systems [31, 11, 34], these series are meticulously collected at specific sampling frequencies. Like musical notes in a composition, they harmonize, revealing patterns that resonate through the symphony of temporal evolution. In this realm, Long-term Time Series Forecasting (LTSF) stands as one of the enduring focal points within the machine learning community, achieving widespread recognition in real-world applications, such as weather forecasting, financial risk assessment, and traffic prediction [23, 19, 25, 22].
@@
Building on the success of various deep LTSF models [41, 45, 43, 42, 24, 19], which primarily leverage neural networks to learn temporal dependencies from discretely sampled data. Currently, researchers [38, 27] have been investigating the application of Koopman theory [40] in recovering continuous dynamic systems, which applies linear evolution operators to analyze dynamic system characteristics in a sufficiently high-dimensional Koopman function space. Nevertheless, the existence of Koopman space relies on the deterministic system, posing challenges given the chaotic nature of real-world time series data, evidenced through the Maximal Lyapunov Exponent in Appendix E.2.
@@
In this paper, inspired by chaos theory [8], we revisit LSTF tasks from a chaos perspective: Linear or complex nonlinear dynamic systems exhibit stable patterns in their trajectories after sufficient evolution, known as attractors. As illustrated in Figure 1(a), attractors can be classified into four types: Fixed Point, indicating stable, invariant systems; Limited cycle, representing periodic behavior; Limited Toroidal, exhibiting quasi-periodic behavior with non-intersecting rings in a 2D plane, reflecting temporal distribution shifts; and Strange Attractor, characterized by nonlinear behavior and complex, non-topological shapes. Supported by chaos theory, we can transcend the limitations of deterministic dynamical systems to construct generalized dynamical system models. Figure 1(b-c) showcases various classical chaotic dynamic systems and the dynamic trajectories of real-world LTSF datasets using the phase space reconstruction method [9]. Notably, the dynamic system trajectories in these LTSF datasets exhibit fixed structures akin to those in typical chaotic systems.
@@
Given this chaos perspective, we consider real-world time series as stemming from an unidentified high-dimensional underlying chaotic system, broadly encompassing nonlinear behaviors beyond periodicity. Our focus centers on recovering continuous chaotic dynamic systems from discretely sampled data for LTSF tasks, with the goal of predicting future time steps through the lens of attractor invariance. Specifically, this problem can be decomposed into two key questions: (i) how to model the underlying continuous chaotic dynamic system based on discretely sampled time series data; (ii) how to enhance forecasting performance by utilizing the attractors within the system.
@@
Figure 1: (a): Classical chaotic systems with noise. (b): Dynamic system structure of real-world datasets. (c): Different types of Attractors. See more figures in Appendix E.1.
@@
In this context, Attraos comes into existence with the goal of capturing the underlying order within the seeming chaos via attractors. For tackling the first question, we employ a non-parametric phase space reconstruction method to recover the temporal dynamics and propose a Multi-resolution Dynamic Memory Unit (MDMU) to memorize the structural dynamics within historical sampled data. Specifically, as polynomials have been proven to be universal approximators for dynamic systems [4], MDMU expands upon the work of the State Space Model (SSM) [12, 10, 18] to different orthogonal polynomial subspaces. This allows for memorizing diverse dynamic structures that encompass various attractor patterns, while theoretically minimizing the boundary of attractor evolution error.
@@
To address the second question, we devise a frequency-enhanced local evolution strategy, which is built upon the recognition that attractor differences are amplified in the frequency domain, as observed in the field of neuroscience [5, 14, 6]. Concretely, for dynamic system components that belong to the same attractor, we apply a consistent evolution operator to derive their future states in the frequency domain. Our contributions can be summarized as follows:
@@
  • Empirical Evidence. Various experiments validate the superior performance of Attraos. Besides Leveraging the properties of chaotic dynamic systems, we explore their extended applications in LTSF tasks, focusing on chaotic evolution, modulation, representation, and reconstruction.
@@ 2 Preliminary
Attractor in Chaos Theory.  In chaos theory, the interaction of three or more variables exhibiting periodic behavior gives rise to a complex dynamic system characterized by chaos. According to Takens’s theorem [29], assuming an ideal dynamic system $\mathcal{F}:\mathcal{M}\to\mathcal{M}$ that “lives” on attractor $\mathcal{A}$ in manifold space $\mathcal{M}$ which locally $\mathcal{C}^{N}$ (N-times differentiable), time series data $\{z_{i}\}\in\mathbb{R}$ can be interpreted as the observation of it by an unknown observation function $h$. To explore the properties of the unknown ideal dynamic system, we can employ the phase space reconstruction (PSR) method to establish an approximation $\mathcal{K}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ which lives in differential homomorphism attractor $\tilde{\mathcal{A}}$ in the Euclidean space with suitable dimension $m$ [9]. The whole process is illustrated in Equation (1), where $\{z_{i}\}$, $\{u_{i}\}$ are the sampled data from two dynamic systems. Strictly speaking, in our paper, the chaotic attractor structure $\tilde{\mathcal{A}}=\{\tilde{\mathcal{A}}_{i}\}$ we focused on is in phase space $\mathbb{R}^{m}$. To facilitate understanding, we further provide a visual example of the Lorenz96 system in Appendix E.3.
In the forecasting stage, the local prediction method emerges as a prominent one: $u_{i+1}=\mathcal{K}^{(i)}(u_{i})$, where the local evolution $\mathcal{K}^{(i)}$ can be either linear or nonlinear neural network [37, 2, 35], with the parameter being shared among the points in the neighborhood of $u_{i}$ or belong to the same local attractor. Considering the universal approximation capabilities of polynomials for dynamic systems, we leverage the polynomial to describe the chaotic dynamic structures.
$\begin{array}[]{ccc}{a}_{i}\in\mathcal{A}\subset\mathcal{M}&\stackrel{{\scriptstyle\mathcal{F}}}{{\longmapsto}}&{a}_{i+1}\in\mathcal{A}\subset\mathcal{M}\\ \downarrow_{h}&&\downarrow_{h}\\ z_{i}\in\mathbb{R}&&z_{i+1}\in\mathbb{R}\\ \downarrow_{PSR}&&\downarrow_{PSR}\\ u_{i}\in\tilde{\mathcal{A}}\subset\mathbb{R}^{m}&\stackrel{{\scriptstyle\mathcal{K}}}{{\longmapsto}}&u_{i+1}\in\tilde{\mathcal{A}}\subset\mathbb{R}^{m}\end{array}$ (1)
$x^{\prime}(t)$ $=\bm{A}x(t)+\bm{B}u(t)$ (2a)
$x(t)$ $=(K\ast u)(t)$ (2b)
$K(t)$ $=e^{t\bm{A}}\bm{B}$ (2c)
@@
$\langle u,\phi_{n}\rangle_{\mu}=\int_{-\infty}^{t}u(s)\phi_{n}(t,s)\omega(t,s)\mathrm{d}s$ (3)
$x_{n}(t)=\int u(s)K_{n}(t,s)\mathbbm{I}(t,s)\mathrm{d}s.$ (4)
When the basis and measure are solely dependent on time $t$, it can be expressed in a convolution form (2b). In this paper, we will utilize this property to project the dynamic trajectories $\{u_{n}\}$ in phase space onto the polynomial spectral domain with kernel $e^{t\bm{A}}\bm{B}$ (2c) for characterization.
@@ 3 Theoretical Analysis & Methods
Figure 2: Overall architecture of Attraos. Initially, the PSR technique is employed to restore the underlying dynamic structures from historical data $\{z_{i}\}$. Subsequently, the dynamic system trajectory is fed into MDMU, projected onto polynomial space $\mathcal{G}_{\theta}^{N}$ using a time window $\theta$ and polynomial order $N$. Gradually, a hierarchical projection is performed to obtain more macroscopic memories of the dynamic system structure. Finally, local evolution operator $\mathcal{K}^{(i)}$ in the frequency domain is employed to obtain future state, thereby for the prediction.
The overall structure of Attraos is depicted in Figure 2. In this section, we will comprehensively describe the components of Attraos, along with the utilization of efficient computational methods.
@@ 3.1 Phase Space Reconstruction
According to chaos theory, the initial step involves constructing a topologically equivalent dynamic structure through the PSR. The preferred embedding method is typically the Coordinate Delay Reconstruction [29], which does not rely on any prior knowledge of the underlying dynamic system. By utilizing the discrete sampled data $\{z_{i}\}$ and incorporating two hyperparameters, namely, embedding dimension $m$, time delay $\tau$, a trajectory $\{u_{i}\}$ in phase space can be constructed by Equation 5.
@@
$\mathcal{K}_{patch}=\operatorname{Unfold}~{}(\mathcal{K},p,p).$ (6)
For multivariate time series data with $C$ variables, we have observed considerable variations in the Lyapunov exponents of each variable, hence a channel-independent strategy [28] is employed to construct a unified dynamic system $\mathcal{K}\subset\mathbb{R}^{m}$. To accelerate model convergence and reduce complexity, we apply non-overlapping patching to obtain $\mathcal{K}_{patch}$ (6). We denote the number of patches as $L$, using $u\in\mathbb{R}^{B\times L\times D}$ to represent the tensor used for computing, where $D=mp$. The determination of $m$ and $\tau$ is achieved by applying the CC method [20] as shown in Appendix C.1.
@@
3.2 Dynamic Representation by MDMU
@@
$\bm{A}=\mathsf{diag}\{-1,-1,\dots\}$ is a rough approximation of normal Hippo-LegT [11] matrix.
@@
Adhere to Mamba [10], we generate $\bm{B}$ and measure window $\theta$ by linear layer, and propose a novel parameterized method (Proposition 1) for $\bm{A}$ to instantiate Equation 2a.
$\bm{A}=\mathsf{Broadcast}_{D}(\mathsf{diag}\{-1,-1,\dots\}),\quad\quad\bm{B}=\mathsf{Linear}_{\bm{B}}(u),\quad\quad\bm{\Delta}/\theta=\mathsf{softplus}(\mathsf{Linear}_{\bm{\Delta}}(u)).$ (7)
where $\bm{A}\in\mathbb{R}^{D\times N}$ represents the dynamic characteristics of the system’s forward evolution in the polynomial space. Due to its diagonal nature, its representational capacity is comparable to $\mathbb{R}^{D\times N\times N}$; Matrix $\bm{B}\in\mathbb{R}^{B\times L\times N}$ controls the process of projecting $u$ onto the polynomial domain like a gate mechanism; The learnable approximation window $\bm{\Delta}\in\mathbb{R}^{B\times L\times D}$, similar to an attention mechanism, enables adaptive focus on specific dynamic structures (attractors).
@@
As shown in Figure 3(a), in practical computations, we need to discretize Equation 2a to fit the discrete dynamical trajectories. We apply zero-order hold (ZOH) discretization [11] to matrix $\bm{A}$, while opting for a combination of Forward Euler discretization for $\bm{B}$ (instead of the commonly used $\overline{\bm{B}}=(\Delta\bm{A})^{-1}(\exp(\Delta\bm{A})-\bm{I})\cdot\Delta\bm{B}$ in SSMs), resulting in a more concise representation.
@@
Figure 3: (a) Discretization of continuous polynomial approximation for sequence data. $g$ represents the optimal polynomial constructed from polynomial bases. (b) MDMU projects the dynamical structure onto different orthogonal subspaces $\mathcal{G}$ and $\mathcal{S}$. (c) Sequential computation for Equation 2a in $\mathcal{O}(L)$ time complexity. (d) Blelloch tree scanning for Equation 2a in $\mathcal{O}(logL)$ by storing intermediate results.
Next, we can project the dynamics trajectory $u$ onto the polynomial domain using the discretized kernel to obtain dynamic representation $x\in\mathbb{R}^{B\times L\times D\times N}$ (9). This process can be achieved with $\mathcal{O}(L)$ complexity by employing sequential computation (Figure 3(c)) or by utilizing the Blelloch scan (Figure 3(d)) to store intermediate results with $\mathcal{O}(logL)$ complexity.
@@
Up to this point, the construction of the underlying continuous dynamic system $\mathcal{K}$ and its application to discretely sampled data ${z_{n}}$ have been established. However, in this scenario, we are still limited to a single representation of the dynamic structures with measure window $\theta$. The strange attractors, on the other hand, are often composed of multiple fundamental topological structures. Therefore, we require a multi-scale hierarchical representation of dynamic structures to capture their complexity.
To address this, as illustrated in Figure 3(b), we progressively increase the length of the window $\theta$ by powers of 2. The region previously approximated by $g^{\theta_{1}}\in\mathcal{G}_{\theta}^{N}$ (left half) and $g^{\theta_{2}}\in\mathcal{G}_{\theta}^{N}$ (right half) will now be approximated by $g^{2\theta}\in\mathcal{G}_{2\theta}^{N}$. Since $\mathcal{G}_{2\theta}$ is the subspace of $\mathcal{G}_{\theta}$, the projection coefficients on these two spaces can be mutually transformed using the linear projection matrix $H$ and its inverse matrix $H^{\dagger}$ based on the odd or even positions along the $L$ dimension in $x$.
@@
Although $\bm{\Delta}$ is akin to a attention mechanism leads to different measure windows, ie., $\theta_{1}\neq\theta_{2}$, it still maintains the linear projection property for up and down projection:$\mathcal{G}_{\theta_{1}},\mathcal{G}_{\theta_{2}}\leftrightarrow\mathcal{G}_{\theta_{1}+\theta_{2}}$. In our illustration, we have used a unified measure window for simplicity.
Iteratively repeating this process enables us to model the dynamic structure from a more macroscopic perspective. Theorem 2 indicates that under this approach, the polynomial projection error has convergence of order $N$. Hyper-parameter analysis can be found in the experiment section.
Remark 3.5.
When the weight function is uniformly equal across the dynamic structure, $H^{\theta_{1}}$ and $H^{\theta_{2}}$ are shared in each projection level as the projection matrix for the left and right interval.
@@
Based on this hierarchical projection, we additionally want to uphold the constancy of attractor patterns throughout the evolution, which is equivalent to minimizing attractor evolution errors. According to Theorem 3, it is imperative to ensure the separation between attractors, denoted as $\nabla_{i}:=\min_{j,j\neq i}\left(\tilde{\mathcal{A}}_{i}^{T}\tilde{\mathcal{A}}_{i}-\tilde{\mathcal{A}}_{i}^{T}\tilde{\mathcal{A}}_{j}\right)$, is sufficiently large. While there is an intersection between the $\mathcal{G}_{\theta}^{N}$ and $\mathcal{G}_{2\theta}^{N}$ spaces, which limits the attainment of a sufficiently large $\nabla$. To address this issue, we define an orthogonal complement space as $\mathcal{G}_{\theta}^{N}=\mathcal{S}_{\theta}^{N}\bigoplus\mathcal{G}_{2\theta}^{N}$, to establish a series of orthogonal function spaces $\{\mathcal{S}_{\theta}^{N},\mathcal{S}_{2\theta}^{N},...,\mathcal{S}_{2^{L}\theta}^{N},\mathcal{G}_{2^{L}\theta}^{N}\}$. We can extend Equation 10 as: (Details in Appendix A.4)
@@
Theorem 3.
Theorem 4 states that the coarsest-grained $\mathcal{G}$ space, along with a series of orthogonal complement $\mathcal{S}$ spaces, can approximate any given dynamic system structure with finite error bounds in Theorem 2.
@@
The hierarchical projection can be implemented explicitly using iterative display or can be efficiently computed through an implicit implementation proposed in Section 3.3.

### v7

@@ Abstract
In long-term time series forecasting (LTSF) tasks, an increasing number of works have acknowledged that discrete time series originate from continuous dynamic systems and have attempted to model their underlying dynamics. Recognizing the chaotic nature of real-world data, our model, Attraos, incorporates chaos theory into LTSF, perceiving real-world time series as low-dimensional observations from unknown high-dimensional chaotic dynamical systems. Under the concept of attractor invariance, Attraos utilizes non-parametric Phase Space Reconstruction embedding along with a novel multi-resolution dynamic memory unit to memorize historical dynamical structures, and evolves by a frequency-enhanced local evolution strategy. Detailed theoretical analysis and abundant empirical evidence consistently show that Attraos outperforms various LTSF methods on mainstream LTSF datasets and chaotic datasets with only one-twelfth of the parameters compared to PatchTST. Code is available at https://github.com/CityMind-Lab/NeurIPS24-Attraos.
@@ 1 Introduction
In the intricate dance of time, time series unfold. Emerged from continuous dynamical systems [36, 11, 39] in the physical world, these series are meticulously collected at specific sampling frequencies. Like musical notes in a composition, they harmonize, revealing patterns that resonate through the symphony of temporal evolution. In this realm, Long-term Time Series Forecasting (LTSF) stands as one of the enduring focal points within the machine learning community, achieving widespread recognition in real-world applications, such as weather forecasting, financial risk assessment, and traffic prediction [28, 22, 30, 27, 18].
@@
Building on the success of various deep LTSF models [49, 54, 52, 50, 29, 22], which primarily leverage neural networks to learn temporal dependencies from discretely sampled data. Currently, researchers [46, 32] have been investigating the application of Koopman theory [48, 25] in recovering continuous dynamical systems, which applies linear evolution operators to analyze dynamical system characteristics in a sufficiently high-dimensional Koopman function space. Nevertheless, the existence of Koopman space relies on the deterministic system, posing challenges given the chaotic nature of real-world time series data, evidenced through the Maximal Lyapunov Exponent in Appendix E.2.
@@
In this paper, inspired by chaos theory [8], we revisit LSTF tasks from a chaos perspective: Linear or complex nonlinear dynamical systems exhibit stable patterns in their trajectories after sufficient evolution, known as attractors. As illustrated in Figure 1(a), attractors can be classified into four types: Fixed Point, indicating stable, invariant systems; Limited cycle, representing periodic behavior; Limited Toroidal, exhibiting quasi-periodic behavior with non-intersecting rings in a 2D plane, reflecting temporal distribution shifts; and Strange Attractor, characterized by nonlinear behavior and complex, non-topological shapes. Supported by chaos theory, we can transcend the limitations of deterministic dynamical systems to construct generalized dynamical system models. Figure 1(b-c) showcases various classical chaotic dynamical systems and the dynamical trajectories of real-world LTSF datasets using the phase space reconstruction method [9]. Notably, the dynamical system trajectories in these LTSF datasets exhibit fixed structures akin to those in typical chaotic systems.
@@
Given this chaos perspective, we consider real-world time series as stemming from an unidentified high-dimensional underlying chaotic system, broadly encompassing nonlinear behaviors beyond periodicity. Our focus centers on recovering continuous chaotic dynamical systems from discretely sampled data for LTSF tasks, with the goal of predicting future time steps through the lens of attractor invariance. Specifically, this problem can be decomposed into two key questions: (i) how to model the underlying continuous chaotic dynamical system based on discretely sampled time series data; (ii) how to enhance forecasting performance by utilizing the attractors within the system.
@@
Figure 1: (a): Classical chaotic systems with noise. (b): dynamical system structure of real-world datasets. (c): Different types of Attractors. See more figures in Appendix E.1.
@@
In this context, Attraos emerges with the goal of capturing the underlying order within the seeming chaos via attractors. For tackling the first question, we employ a non-parametric phase space reconstruction method to recover the temporal dynamics and propose a Multi-resolution Dynamic Memory Unit (MDMU) to memorize the structural dynamics within historical sampled data. Specifically, as polynomials have been proven to be universal approximators for dynamical systems [4], MDMU expands upon the work of the State Space Model (SSM) [12, 10, 18] to different orthogonal polynomial subspaces. This allows for memorizing diverse dynamical structures that encompass various attractor patterns, while theoretically minimizing the boundary of attractor evolution error.
@@
To address the second question, we devise a frequency-enhanced local evolution strategy, which is built upon the recognition that attractor differences are amplified in the frequency domain, as observed in the field of neuroscience [5, 14, 6]. Concretely, for dynamical system components that belong to the same attractor, we apply a consistent evolution operator to derive their future states in the frequency domain. Our contributions can be summarized as follows:
@@
  • Empirical Evidence. Various experiments validate the superior performance of Attraos. Besides Leveraging the properties of chaotic dynamical systems, we explore their extended applications in LTSF tasks, focusing on chaotic evolution, modulation, representation, and reconstruction.
@@ 2 Preliminary
Attractor in Chaos Theory.  In chaos theory, the interaction of three or more variables exhibiting periodic behavior gives rise to a complex dynamical system characterized by chaos. According to Takens’s theorem [43, 34], assuming an ideal dynamical system $\mathcal{F}:\mathcal{M}\to\mathcal{M}$ that “lives” on attractor $\mathcal{A}$ in manifold space $\mathcal{M}$ which locally $\mathcal{C}^{N}$ (N-times differentiable), time series data $\{z_{i}\}\in\mathbb{R}$ can be interpreted as the observation of it by an unknown observation function $h$. To explore the properties of the unknown ideal dynamical system, we can employ the phase space reconstruction (PSR) method to establish an approximation $\mathcal{K}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ which lives in differential homomorphism attractor $\tilde{\mathcal{A}}$ in the Euclidean space with suitable dimension $m$ [9]. The whole process is illustrated in Equation (1), where $\{z_{i}\}$, $\{u_{i}\}$ are the sampled data from two dynamical systems. Strictly speaking, in our paper, the chaotic attractor structure $\tilde{\mathcal{A}}=\{\tilde{\mathcal{A}}_{i}\}$ we focused on is in phase space $\mathbb{R}^{m}$. To facilitate understanding, we further provide a visual example of the Lorenz96 system in Appendix E.3.
In the forecasting stage, the local prediction method emerges as a prominent one: $u_{i+1}=\mathcal{K}^{(i)}(u_{i})$, where the local evolution $\mathcal{K}^{(i)}$ can be either linear or nonlinear neural network [45, 2, 40], with the parameter being shared among the points in the neighborhood of $u_{i}$ or belong to the same local attractor. Considering the universal approximation capabilities of polynomials for dynamical systems, we leverage the polynomial to describe the chaotic dynamical structures.
$\begin{array}[]{ccc}{a}_{i}\in\mathcal{A}\subset\mathcal{M}&\stackrel{{\scriptstyle\mathcal{F}}}{{\longmapsto}}&{a}_{i+1}\in\mathcal{A}\subset\mathcal{M}\\ \downarrow_{h}&&\downarrow_{h}\\ z_{i}\in\mathbb{R}&&z_{i+1}\in\mathbb{R}\\ \downarrow_{PSR}&&\downarrow_{PSR}\\ u_{i}\in\tilde{\mathcal{A}}\subset\mathbb{R}^{m}&\stackrel{{\scriptstyle\mathcal{K}}}{{\longmapsto}}&u_{i+1}\in\tilde{\mathcal{A}}\subset\mathbb{R}^{m}\end{array}$ (1) $x^{\prime}(t)$ $=\bm{A}x(t)+\bm{B}u(t)$ (2a) $x(t)$ $=(K\ast u)(t)$ (2b) $K(t)$ $=e^{t\bm{A}}\bm{B}$ (2c)
@@
$\langle u,\phi_{n}\rangle_{\mu}=\int_{-\infty}^{t}u(s)\phi_{n}(t,s)\omega(t,s)\mathrm{d}s$ (3) $x_{n}(t)=\int u(s)K_{n}(t,s)\mathbbm{I}(t,s)\mathrm{d}s$ (4)
When the basis and measure are solely dependent on time $t$, it can be expressed in a convolution form (2b). In this paper, we will utilize this property to project the dynamical trajectories $\{u_{n}\}$ in phase space onto the polynomial spectral domain with kernel $e^{t\bm{A}}\bm{B}$ (2c) for characterization.
@@ 3 Theoretical Analysis & Methods
Figure 2: Overall architecture of Attraos. Initially, the PSR technique is employed to restore the underlying dynamical structures from historical data $\{z_{i}\}$. Subsequently, the dynamical system trajectory is fed into MDMU, projected onto polynomial space $\mathcal{G}_{\theta}^{N}$ using a time window $\theta$ and polynomial order $N$. Gradually, a hierarchical projection is performed to obtain more macroscopic memories of the dynamical system structure. Finally, local evolution operator $\mathcal{K}^{(i)}$ in the frequency domain is employed to obtain future state, thereby for the prediction.
The overall structure of Attraos is illustrated in Figure 2. In this section, we provide a comprehensive description of its components, including the Phase Space Reconstruction embedding, the Multi-Resolution Dynamic Memory Unit (MDMU), and the frequency-enhanced local evolution, as well as the efficient computational methods employed.
@@ 3.1 Phase Space Reconstruction
According to chaos theory, the initial step involves constructing a topologically equivalent dynamical structure through the PSR. The preferred embedding method is typically the Coordinate Delay Reconstruction [34], which does not rely on any prior knowledge of the underlying dynamical system. By utilizing the discretely sampled data $\{z_{i}\}$ and incorporating two hyper-parameters, namely, embedding dimension $m$ and time delay $\tau$, a high-dimensional dynamical trajectory $\{u_{i}\}$ in phase space can be constructed by Eq. (5).
@@
$\mathcal{K}_{patch}=\operatorname{Unfold}~{}(\mathcal{K},p,p)$ (6)
For multivariate time series data with $C$ variables, we have observed considerable variations in the Lyapunov exponents of each variable, hence a channel-independent strategy [33] is employed to construct a unified dynamical system $\mathcal{K}\subset\mathbb{R}^{m}$. To accelerate model convergence and reduce complexity, we apply non-overlapping patching to obtain $\mathcal{K}_{patch}$ (6). We denote the number of patches as $L$, using $u\in\mathbb{R}^{B\times L\times D}$ to represent the tensor used for computing, where $D=mp$. The determination of $m$ and $\tau$ is achieved by applying the CC method [23] as shown in Appendix C.1.
@@
Remark 3.2.
In a large body of dynamical literature [44, 41, 20, 25], local linear approximation serves as an effective method for modeling dynamical systems, providing a basis for the effectiveness of the patching operations in this paper.
3.2 Dynamical Representation by MDMU
@@
$\bm{A}=\mathsf{diag}\{-1,-1,\dots\}$ is a rough approximation of normal Hippo-LegT [11] matrix, which utilizes polynomial projection under a finite measure window (Lebesgue measure).
@@
Adhere to Mamba [10], we generate $\bm{B}$ and measure window $\theta$ by linear layer, and propose a novel parameterized method (Proposition 1) for $\bm{A}$ to instantiate Eq. (2a):
$\bm{A}=\mathsf{Broadcast}_{D}(\mathsf{diag}\{-1,-1,\dots\}),\quad\quad\bm{B}=\mathsf{Linear}_{\bm{B}}(u),\quad\quad\bm{\Delta}/\theta=\mathsf{softplus}(\mathsf{Linear}_{\bm{\Delta}}(u)),$ (7)
where $\bm{A}\in\mathbb{R}^{D\times N}$ represents the dynamical characteristics of the system’s forward evolution in the polynomial space. Due to its diagonal nature, its representational capacity is comparable to $\mathbb{R}^{D\times N\times N}$; Matrix $\bm{B}\in\mathbb{R}^{B\times L\times N}$ controls the process of projecting $u$ onto the polynomial domain like a gate mechanism; The learnable approximation window $\bm{\Delta}\in\mathbb{R}^{B\times L\times D}$, similar to an attention mechanism, enables adaptive focus on specific dynamical structures (attractors).
@@
As shown in Figure 3(a), in practical computations, we need to discretize Eq. (2a) to fit the discrete dynamical trajectories. We apply zero-order hold (ZOH) discretization [11] to matrix $\bm{A}$, while opting for a combination of Forward Euler discretization for $\bm{B}$ (instead of the commonly used $\overline{\bm{B}}=(\Delta\bm{A})^{-1}(\exp(\Delta\bm{A})-\bm{I})\cdot\Delta\bm{B}$ in SSMs), resulting in a more concise representation:
@@
Figure 3: (a) Discretization of continuous polynomial approximation for sequence data. $g$ represents the optimal polynomial constructed from polynomial bases. (b) MDMU projects the dynamical structure onto different orthogonal subspaces $\mathcal{G}$ and $\mathcal{S}$. (c) Sequential computation for Eq. (2a) in $\mathcal{O}(L)$ time complexity. (d) Blelloch tree scanning for Eq. (2a) in $\mathcal{O}(logL)$ by storing intermediate results.
Next, we can project the dynamical trajectory $u$ onto the polynomial domain using the discretized kernel to obtain dynamical representation $x\in\mathbb{R}^{B\times L\times D\times N}$ (9). This process can be achieved with $\mathcal{O}(L)$ complexity by employing sequential computation (Figure 3(c)) or by utilizing the Blelloch scan (Figure 3(d)) to store intermediate results with $\mathcal{O}(logL)$ complexity.
@@
Up to this point, the construction of the underlying continuous dynamical system $\mathcal{K}$ and its application to discretely sampled data ${z_{n}}$ have been established. However, in this scenario, we are still limited to a single representation of the dynamical structures with measure window $\theta$. The strange attractors, on the other hand, are often composed of multiple fundamental topological structures. Therefore, we require a multi-scale hierarchical representation of dynamical structures to capture their complexity.
To address this, as illustrated in Figure 3(b), we progressively increase the length of the window $\theta$ by powers of 2. The region previously approximated by $g^{\theta_{1}}\in\mathcal{G}_{\theta}^{N}$ (left half) and $g^{\theta_{2}}\in\mathcal{G}_{\theta}^{N}$ (right half) will now be approximated by $g^{2\theta}\in\mathcal{G}_{2\theta}^{N}$.
The projection coefficients on these two spaces can be mutually transformed using the linear projection matrix $H$ and its inverse matrix $H^{\dagger}$ based on the odd or even positions along the $L$ dimension in $x$.
@@
Although $\bm{\Delta}$ is akin to an attention mechanism leads to different measure windows, ie., $\theta_{1}\neq\theta_{2}$, it still maintains the linear projection property for up and down projection:$\mathcal{G}_{\theta_{1}},\mathcal{G}_{\theta_{2}}\leftrightarrow\mathcal{G}_{\theta_{1}+\theta_{2}}$. In our illustration, we have used a unified measure window for simplicity.
@@
Iteratively repeating this process enables us to model the dynamical structure from a more macroscopic perspective. Theorem 2 indicates that under this approach, the polynomial projection error has convergence of order $N$. Hyper-parameter analysis can be found in Figure 5.
Remark 3.6.
When the weight function is uniformly equal across the dynamical structure, $H^{\theta_{1}}$ and $H^{\theta_{2}}$ are shared in each projection level as the projection matrix for the left and right interval.
Theorem 3.
Based on this hierarchical projection, we additionally want to uphold the constancy of attractor patterns throughout the evolution, which is equivalent to minimizing attractor evolution errors. According to Theorem 3, it is imperative to ensure the separation between attractors, denoted as $\nabla_{i}:=\min_{j,j\neq i}\left(\tilde{\mathcal{A}}_{i}^{T}\tilde{\mathcal{A}}_{i}-\tilde{\mathcal{A}}_{i}^{T}\tilde{\mathcal{A}}_{j}\right)$, is sufficiently large. While there is an intersection between the $\mathcal{G}_{\theta}^{N}$ and $\mathcal{G}_{2\theta}^{N}$ spaces, which limits the attainment of a sufficiently large $\nabla$. To address this issue, we define an orthogonal complement space as $\mathcal{G}_{\theta}^{N}=\mathcal{S}_{\theta}^{N}\bigoplus\mathcal{G}_{2\theta}^{N}$, to establish a series of orthogonal function spaces $\{\mathcal{S}_{\theta}^{N},\mathcal{S}_{2\theta}^{N},...,\mathcal{S}_{2^{L}\theta}^{N},\mathcal{G}_{2^{L}\theta}^{N}\}$. We can extend Eq. (12) as:
@@
Theorem 4 states that the coarsest-grained $\mathcal{G}$ space, along with a series of orthogonal complement $\mathcal{S}$ spaces, can approximate any given dynamical system structure with finite error bounds in Theorem 2.
@@
The hierarchical projection can be implemented explicitly using iterative display or can be efficiently computed through an implicit implementation proposed in the following Section 3.3.